# Proca stars in excited states

### v1

@@ IV.1 Families of stationary solutions for ground and excited states
The set of ordinary differential equations is integrated numerically using a fourth-order Runge-Kutta method. The code takes a central value of the potential scalar $\varphi_{0}$ as an input parameter, and then employs a shooting algorithm to determine the value of the frequency $\omega$ that corresponds to the asymptotically decaying solution of the scalar potential. For simplicity, we set the value of the Proca mass parameter to $m=1$, but notice that the solutions can later be rescaled to arbitrary values of $m$ since our system of equations is invariant under the rescaling:
In this section we will present the properties of the stationary configurations for three families of Proca star solutions, corresponding to the ground state and the first two excited states. Table 1 shows the parameters that correspond to the maximum mass configuration for all three families of solutions. In particular we show the central value of the scalar potential $\varphi_{0}$, the corresponding frequency $\omega$, the total mass $M$, the binding energy $U$, as well as the effective radius $R_{99}$ and effective compactness defined as $C_{99}:=M/R_{99}$. Our data can be compared with that already reported in [27, 14, 28] for the ground state. However, we also explore the properties of the first two excited states, which correspond to solutions with one or two nodes in the vector potential $a(r)$ (notice the scalar potential $\varphi(r)$ already has one node even in the ground state, see below).
@@
Figure 1: Left panel: Total integrated mass $M$ (in units of $M^{2}_{Planck}/m$) as a function of frequency $\omega$. We observe that as the excitation level of the star increases the total mass $M$ also increases. Right panel: Binding energy $U$ as function of $\varphi_{0}$. For each of the families $U$ has a global minimum, which corresponds to the solutions with maximum mass of Table 1.
We show our three different families of solutions in Figure 1. The left panel shows the total integrated mass $M$ as a function of the oscillation frequency $\omega$, while the right panel shows the binding energy $U$ as a function of the central value of the scalar potential $\varphi_{0}$. We can observe that excited states have in general higher masses. Additionally, we also find that the frequency domain is bounded for all three families, showing the characteristic spiral pattern in the mass versus frequency plot that also appears for boson stars, with the minimum frequency increasing with the excitation level. Also, for all three families there is a value of the central field for which the binding energy reaches a minimum, which in fact corresponds to the solutions with maximum mass. As we increase the value of $\varphi_{0}$ further the binding energy also increases and eventually becomes positive, corresponding to solutions that are no longer gravitationally bound.
Solutions with negative binding energy in all three families can be further separated into solutions to the left and to the right of the minimum binding energy $U$ (maximum mass $M$) in the right panel of Figure 1. Naively, one could expect that all solutions with negative binding energy should be stable under perturbations since they represent gravitationally bound stars. However, as has already been shown [14], this is not true even for solutions in the ground state: configurations to the right of the minimum are unstable to perturbations and can either collapse to a black hole or migrate to a configuration on the stable left branch (this is the same in the case of boson stars). Having a solution with $U<0$ is therefore a necessary but not sufficient condition for the stability of the Proca stars in the ground state. The purpose of this paper is to investigate if something similar happens for solutions in excited states. We will return to this point when we consider dynamical simulations below.
@@
In Figure 2 we consider the effective radius and compactness of our three families of solutions. The right panel shows the total mass $M$ as a function of the effective radius $R_{99}$, while the left panel shows the effective compactness defined as $C_{99}=M/R_{99}$ as a function of $\varphi_{0}$. Notice that despite the fact that Proca stars in the ground state have in general a lower mass than those in excited states, they exhibit greater compactness. This is not too surprising since Proca stars in excited states are generally larger in size, as can be seen in the right panel of the Figure. For the ground state family the effective compactness reaches a maximum value of $C_{99}\simeq 0.12$. This value is still far from $C=0.5$ corresponding to the compactness for a Schwarzschild black hole, but is comparable with other kind of compact objects [29].
In order to better illustrate the properties of our solutions we consider three particular configurations in Figure 3. We consider three solutions with the same central value of the scalar potential $\varphi_{0}=0.081$: One in the ground state, one in the first excited state, and one in the second excited state. All three configurations can be shown to have negative binding energy. Panel (a) shows the density profile $\rho(r)$ for all three solutions. We notice that the ground state possesses a single core of matter, while the excited states have a core and shells surrounding it, and as a result the effective radii of the excited configurations are larger. The first excited state has one such shell outside the core, while the second excited state has two shells (compare this with the analogue case for excited boson stars in [30]).
Figure 3: We show the solutions for the ground state and the first two excited states for the same central value of the scalar potential $\varphi_{0}=0.081$ (the color scheme is the same for all four figures). Panel (a) shows energy density $\rho(r)$. Panels (b) , (c), and (d) show the profiles for the scalar potential $\varphi(r)$, the electric field $e(r)$, and radial component of vector potential $a(r)$.
The other three panels of the figure show the profile functions $\varphi(r)$, $e(r)$ and $a(r)$ for the ground and first two excited states. All these functions decay rapidly far away, as expected. We can also observe that the number of nodes increases with the excitation levels. The scalar function $\varphi(r)$ in the ground state already has one node, followed by the first and second excited states with two and three nodes, respectively. Additionally, in the ground state both the vector potential $a(r)$ and the Proca electric field $e(r)$ have no nodes, while they have one node in the first excited state and two nodes in the second.
@@ IV.2 Dynamical evolutions and stability
The values presented in Table 1 will serve as a reference for choosing models for our dynamical evolutions of perturbed Proca star configurations. In all the numerical simulations presented below we always add a small Gaussian perturbation centered at the origin and with unit width to the scalar potential $\varphi(r)$, while leaving the vector potential $a(r)$ unchanged. The amplitude of this perturbation is always taken as 5% of the maximum amplitude of the scalar potential of the original stationary solution. As mentioned above, once we add the perturbation to $\varphi(r)$ we solve again the the Hamiltonian and Gauss constraints to obtain self-consistent initial data.
@@ IV.2.1 First excited state
Figure 4: Binding energy $U$ as function of $\varphi_{0}$ for the first excited state. The circles show the 7 models that we considered for our perturbed dynamical evolutions. The color bar indicates the scale of the total integrated mass $M$ of the unperturbed stationary configurations. The size of each circle in the plot is proportional to the total mass, and increases as we approach the the minimum of the binding energy.
Figure 6: Apparent horizon mass a a function of time time for the three models in the first excited state that exhibit collapse to a black hole. The time at which the code first finds an apparent horizon corresponds to the dashed vertical line. The dash-dot horizontal line represents the value of the total integrated mass of the spacetime at $t=0$.
@@
The central lapse collapses to zero for the models with $\varphi_{0}=$ 0.081, 0.038, and 0.022, indicating the formation of a black hole. The behavior of the lapse is similar in all 3 models, the main difference being the time at which the collapse occurs. In particular, we find that if the star’s mass is closer to the maximum (binding energy closer to the minimum) it becomes more unstable, leading to a faster collapse. In order to confirm the formation of a black hole we look for apparent horizons during the evolution. Figure 6 shows the apparent horizon mass as a function of time. We can observe that the apparent horizon suddenly appears in all three cases (at different times), coinciding with the time at which the central lapse collapses to zero. The horizon mass is always smaller that the initial total mass of the spacetime (shown as a blue dashed line). We only show the horizon mass for a short time after its formation, as later the numerical errors become too large due to the well known phenomenon of slice stretching (we are evolving without a shift vector).
Figure 7: Evolution of the real part of the scalar potential $\phi$ at $r=0$ for the models with $\varphi_{0}=0.006$ (upper panel) and $\varphi_{0}=0.038$ (lower panel). These models illustrate the behavior of the field’s oscillations when the Proca star is either migrating to another more stable solution, or collapsing into black hole.
@@
We can analyze the final state of our configurations further by doing a Fast Fourier Transform (FFT) of the real part of the scalar potential $\phi$ evaluated at $r=0$, both at the beginning and at the final stages of the time evolutions. This is shown in Figure 9. The dashed blue line shows the FFT for the initial stages of the evolution for all 4 migrating models. In each case we can see a single peak frequency $\omega_{i}$, which coincides with the oscillating frequency of our original stationary solution. The solid red lines correspond to the FFT for the late stages of our evolutions. The spectra for the different models have clearly changed in time, and now show a very different dominant frequency $\omega_{f}$, plus a few other lower amplitude modulating frequencies. To obtain the FFT’s, we divide the evolution into initial and final stages based on the behavior of the oscillation of the real part of the scalar potential at $r=0$. We select an appropriate time interval for each migrating case. Generally, the initial stage consists of only a short time interval because the oscillations of the perturbed solution increase rapidly. The final stage corresponds to a later time when the oscillations have settled into a regular modulated pattern. We only exclude a short portion of the evolution, which represents the transition phase in the middle of the evolutions.
@@ IV.2.2 Second excited state
Figure 11: Binding energy $U$ as function of $\varphi_{0}$ for the second excited state. The circles show the 6 models that we considered for our perturbed dynamical evolutions. As before, the color bar and the size of the circles indicate the initial total mass.
@@
Figure 12 shows again the central value of the lapse function $\alpha$ as a function of time for these 6 models. Similar behaviors are observed, with migration appearing for the 4 lower mass models, and collapse to a black hole for the other 2. Figure 13 shows the apparent horizon mass as a function of time for the two collapsing models. We can observe again that the apparent horizon suddenly appears at a time coinciding with the moment at which the central lapse collapses to zero. It is however important to mention that in fact the majority of Proca stars configurations in the second excited state experience collapse to a black hole, with only configurations with very low masses potentially presenting signs of migration. The fact that 4 out of our 6 models migrate is because we have selected them on purpose.
@@
Let us now return to the 4 migrating models with $\varphi_{0}=$ 0.0099, 0.0056, 0.0042 and 0.0023. All these models display irregular oscillations in the central value of the lapse function around some equilibrium position. The total integrated mass at the boundary as a function of time for these models is shown in Figure 15. Here, we again observe a fractional loss of the initial mass due to the fact that some of the field is ejected to infinity (in all these cases approximately 35-39%).
@@ V Discussion and conclusions
In this paper we have considered families of solutions for excited states of Proca stars in spherical symmetry, with a focus on exploring the dynamical behavior of the first two excited states under perturbations. We considered both the families of stationary solutions, as well as non-linear dynamical simulations of a set of self-consistent perturbed models. Our findings offer valuable insights into the stability and evolution of excited Proca stars, contributing to the further understanding of these compact objects.
It is well known that Proca stars in the ground state are stable against small perturbations for a specific branch corresponding to those configurations with negative binding energy and mass to the left of the maximum mass (minimum binding energy) in the mass versus $\varphi_{0}$ diagram. However, we find that in the case of Proca stars in the first and second excited states this stable branch does not exist. In fact, we find that for excited states only a metastable branch exists for configurations with negative binding energy and low masses. In this metastable branch the Proca stars migrate under perturbations to solutions with lower mass in the stable branch of the ground state, while ejecting excess field to infinity. Our dynamical simulations have allowed us not only to observe the migration phenomena for this type of configurations, but also to determine the final configuration on the ground state to which they are migrating.
For the case of excited states, the absence of dynamically stable Proca stars raises intriguing questions about the factors influencing their instability. Notably, our model does not incorporate self-interactions, which have been demonstrated to play a crucial role in stabilizing excited boson stars [30] and fermion-boson stars [31]. Introducing self-interaction parameters into the model presents an opportunity to further explore the stability and dynamics of excited Proca stars. We are currently in the process of studying this further.

### v2

@@ IV Results
Before discussing our main results, we will say something about the numerical methods used here. For finding the stationary solutions, as well as initial data for the perturbed solutions, the system of ordinary differential equations is integrated numerically using a fourth-order Runge-Kutta method. The code takes a central value of the potential scalar $\varphi_{0}$ as an input parameter, and employs a shooting algorithm to determine the value of the frequency $\omega$ that corresponds to the asymptotically decaying solution of the scalar potential. For simplicity, we set the value of the Proca mass parameter to $m=1$, but notice that the solutions can later be rescaled to arbitrary values of $m$ since our system of equations is invariant under the rescaling:
Although we do not show them here, we have performed extensive tests to check that our numerical solutions do indeed converge to fourth order, as expected. In particular, we have checked that the Hamiltonian and Gauss constraints converge to zero. Typical solutions are obtained with a spatial resolution of $\Delta r=0.01$, with the boundaries located at $r\simeq 100$. With this resolution, the maximum violations of the Hamiltonian and Gauss constraints are of the order of $10^{-9}$.
For a given value of $\varphi_{0}$, and a given resolution, we determine the frequency $\omega$ to machine round-off error. However, the value of $\omega$ does change slightly with resolution. For example, when changing the resolution from $\Delta r=0.01$ to $\Delta r=0.02$, while keeping the same value of $\varphi_{0}$ and the same position of the boundaries, we typically find changes in the value of the frequency $\omega$ in the 11th significant figure.
As mentioned above, for the dynamical evolutions we use the OllinSphere code. This is a general purpose finite difference numerical relativity code in spherical symmetry based on the BSSN formulation, that uses fourth-order finite differencing in space, and a method of lines evolution in time with a fourth-order Runge–Kutta integrator. This code has been extensively tested for consistency and accuracy in many different scenarios [38, 39, 40, 41, 42].
For our dynamical simulations we typically use a larger value of the spatial increment, $\Delta r=0.05$, with a Courant parameter $\Delta t/\Delta r=0.5$ to ensure numerical stability. This is to reduce computational demands since the evolutions can be extremely long, with several million time steps, in order to properly see the late-time behavior. We monitor the constraint violation during our simulations, and typically find that the root mean square (RMS) norm of the constraints grows slowly with time, but converges to zero with increased resolution. We have also checked that moving the boundaries further away does not affect our results in a significant way.
@@ IV.1 Families of stationary solutions for ground and excited states
In this section we will present the properties of the stationary configurations for three families of Proca star solutions, corresponding to the ground state and the first two excited states. Table 1 shows the parameters that correspond to the maximum mass configuration for all three families. In particular, we show the central value of the scalar potential $\varphi_{0}$, the corresponding frequency $\omega$, the total mass $M$, the binding energy $U$, as well as the effective radius $R_{99}$ and effective compactness defined as $C_{99}:=M/R_{99}$. Our data can be compared with that reported in [43, 18, 44] for the ground state. However, we also explore the properties of the first two excited states, which correspond to solutions with one or two nodes in the vector potential $a(r)$ (notice the scalar potential $\varphi(r)$ already has one node even in the ground state, see below).
@@
Figure 1: Left panel: Total integrated mass $M$ (in units of $M^{2}_{Planck}/m$) as a function of frequency $\omega$. We observe that as the excitation level of the star increases the total mass $M$ also increases. Right panel: Binding energy $U$ as a function of $\varphi_{0}$. For each of the families $U$ has a global minimum, which corresponds to the solutions with maximum mass of Table 1.
We show our three different families of solutions in Figure 1. The left panel shows the total integrated mass $M$ as a function of the oscillation frequency $\omega$, while the right panel shows the binding energy $U$ as a function of the central value of the scalar potential $\varphi_{0}$. We can observe that excited states have, in general, higher masses. Additionally, we also find that the frequency domain is bounded for all three families, showing the characteristic spiral pattern in the mass versus frequency plot that also appears for boson stars, with the minimum frequency increasing with the excitation level. Also, for all three families, there is a value of the central field for which the binding energy reaches a minimum, which in fact corresponds to the solutions with maximum mass. As we increase the value of $\varphi_{0}$ further, the binding energy also increases and eventually becomes positive, corresponding to solutions that are no longer gravitationally bound.
Solutions with negative binding energy in all three families can be further separated into solutions to the left and to the right of the minimum binding energy $U$ (maximum mass $M$) in the right panel of Figure 1. Naively, one could expect that all solutions with negative binding energy should be stable under perturbations since they represent gravitationally bound stars. However, as has already been shown [18], this is not true even for solutions in the ground state: configurations to the right of the minimum in the $U$ vs. $\phi_{0}$ plot are unstable to perturbations, and can either collapse to a black hole or migrate to a configuration on the stable left branch (this is the same in the case of boson stars). Having a solution with $U<0$ is therefore a necessary but not sufficient condition for the stability of the Proca stars in the ground state. The purpose of this paper is to investigate whether something similar happens for solutions in excited states. We will return to this point when we consider dynamical simulations below.
@@
In Figure 2 we consider the effective radius and compactness of our three families of solutions. The right panel shows the total mass $M$ as a function of the effective radius $R_{99}$, while the left panel shows the effective compactness defined as $C_{99}=M/R_{99}$ as a function of $\varphi_{0}$. Notice that despite the fact that Proca stars in the ground state have in general a lower mass than those in excited states, they exhibit greater compactness. This is not too surprising, since Proca stars in excited states are generally larger in size, as can be seen in the right panel of the figure. For the ground state family the effective compactness reaches a maximum value of $C_{99}\simeq 0.12$. This value is still far from $C=0.5$ corresponding to the compactness of a Schwarzschild black hole, but is comparable with other kind of compact objects [45].
In order to better illustrate the properties of our solutions, we consider three particular configurations in Figure 3. We choose three solutions with the same central value of the scalar potential $\varphi_{0}=0.081$: One in the ground state, one in the first excited state, and one in the second excited state. All three configurations can be shown to have negative binding energy. Panel (a) shows the density profile $\rho(r)$ for all three solutions. We notice that the ground state possesses a single core of matter, while the excited states have a core and shells surrounding it, and as a result the effective radii of the excited configurations are larger. The first excited state has one such shell outside the core, while the second excited state has two shells (compare this with the analogue case for excited boson stars in [24]).
Figure 3: We show the solutions for the ground state and the first two excited states for the same central value of the scalar potential $\varphi_{0}=0.081$ (the color scheme is the same for all four figures). Panel (a) shows energy density $\rho(r)$. Panels (b), (c), and (d) show the profiles for the scalar potential $\varphi(r)$, the electric field $e(r)$, and the radial component of vector potential $a(r)$.
The other three panels of the figure show the profile functions $\varphi(r)$, $e(r)$ and $a(r)$ for the ground state and the first two excited states. All these functions decay rapidly far away, as expected. We can also observe that the number of nodes increases with the excitation levels. The scalar function $\varphi(r)$ in the ground state already has one node, followed by the first and second excited states with two and three nodes, respectively. Additionally, in the ground state both the vector potential $a(r)$ and the Proca electric field $e(r)$ have no nodes, while they have one node in the first excited state and two nodes in the second.
@@ IV.2 Dynamical evolutions and stability
The values presented in Table 1 will serve as a reference for choosing models for our dynamical evolutions of perturbed Proca star configurations. In all the numerical simulations presented below, we always add a small Gaussian perturbation centered at the origin and with unit width to the scalar potential $\varphi(r)$, while leaving the vector potential $a(r)$ unchanged. The amplitude of this perturbation is always taken as 5% of the maximum amplitude of the scalar potential of the original stationary solution. As mentioned above, once we add the perturbation to $\varphi(r)$ we solve again the Hamiltonian and Gauss constraints to obtain self-consistent initial data.
We should mention here the fact that we have settled on a 5% perturbation in order to both be consistent across all simulations, and also to have a large enough perturbation that allows us to see the final stages of the evolution in a reasonable computational time, be it collapse to a black hole, migration, or dispersion. We have checked that smaller perturbations give similar results, but the instabilities take much longer to develop. Notice that for unstable configurations one does not need to add a perturbation by hand in order to trigger an instability, since numerical truncation error will trigger it on its own given enough time (we have checked that this is indeed the case). However, this has two disadvantages: First, for high resolutions the numerical truncation error is very small and the instability can take a very long time to develop, and second, triggering a numerical instability with pure truncation error makes it impossible to check for convergence of the late time behavior since at higher resolution the effective perturbation is smaller. Because of this, we prefer to add a finite self-consistent perturbation to our stationary solutions as described above.
@@ IV.2.1 First excited state
Figure 4: Binding energy $U$ as a function of $\varphi_{0}$ for the first excited state. The circles show the 7 models that we considered for our perturbed dynamical evolutions. The color bar indicates the scale of the total integrated mass $M$ of the unperturbed stationary configurations. The size of each circle in the plot is proportional to the total mass, and increases as we approach the the minimum of the binding energy.
The central lapse collapses to zero for the models with $\varphi_{0}=$ 0.081, 0.038, and 0.022, indicating the formation of a black hole. The behavior of the lapse is similar in all 3 models, the main difference being the time at which the collapse occurs. In particular, we find that if the star’s mass is closer to the maximum (binding energy closer to the minimum) it becomes more unstable, leading to a faster collapse. In order to confirm the formation of a black hole we look for apparent horizons during the evolution. Figure 6 shows the apparent horizon mass as a function of time. We can observe that the apparent horizon suddenly appears in all three cases (at different times), coinciding with the time at which the central lapse collapses to zero. The horizon mass is always smaller than the initial total mass of the spacetime (shown as a blue dashed line). We only show the horizon mass for a short time after its formation, as later the numerical errors become too large due to the well known phenomenon of slice stretching (we are evolving without a shift vector).
Figure 6: Apparent horizon mass as a function of time time for the three models in the first excited state that exhibit collapse to a black hole. The time at which the code first finds an apparent horizon corresponds to the dashed vertical line. The dash-dot horizontal line represents the value of the total integrated mass of the spacetime at $t=0$.
Figure 7: Evolution of the real part of the scalar potential $\phi$ at $r=0$ for the models with $\varphi_{0}=0.006$ (upper panel) and $\varphi_{0}=0.038$ (lower panel). These models illustrate the behavior of the field’s oscillations when the Proca star is either migrating to another more stable solution, or collapsing into a black hole.
@@
We can analyze the final state of our configurations further by doing a Fast Fourier Transform (FFT) of the real part of the scalar potential $\phi$ evaluated at $r=0$, both at the beginning and at the final stages of the time evolutions. This is shown in Figure 9. The dashed blue line shows the FFT for the initial stages of the evolution for all 4 migrating models. In each case, we can see a single peak frequency $\omega_{i}$, which coincides with the oscillating frequency of our original stationary solution. The solid red lines correspond to the FFT for the late stages of our evolutions. The spectra for the different models have clearly changed in time, and now show a very different dominant frequency $\omega_{f}$, plus a few other lower amplitude modulating frequencies (we have checked that when doing longer simulations the height of those secondary peaks becomes smaller, showing that those modulating oscillations decay away slowly). To obtain the FFT’s, we divide the evolution into initial and final stages based on the behavior of the oscillation of the real part of the scalar potential at $r=0$. We select an appropriate time interval for each migrating case. Generally, the initial stage consists of only a short time interval because the oscillations of the perturbed solution increase rapidly. The final stage corresponds to a later time when the oscillations have settled into a regular modulated pattern. We only exclude a short portion of the evolution, which represents the transition phase in the middle of the evolutions.
@@ IV.2.2 Second excited state
Figure 11: Binding energy $U$ as a function of $\varphi_{0}$ for the second excited state. The circles show the 6 models that we considered for our perturbed dynamical evolutions. As before, the color bar and the size of the circles indicate the initial total mass.
@@
Figure 12 shows again the central value of the lapse function $\alpha$ as a function of time for these 6 models. Similar behaviors are observed, with migration appearing for the 4 lower mass models, and collapse to a black hole for the other 2. Figure 13 shows the apparent horizon mass as a function of time for the two collapsing models. We can observe again that the apparent horizon suddenly appears at a time coinciding with the moment at which the central lapse collapses to zero. It is, however, important to mention that in fact the majority of Proca star configurations in the second excited state experience collapse to a black hole, with only configurations with very low masses potentially presenting signs of migration. The fact that 4 out of our 6 models migrate is because we have selected them on purpose.
@@
Let us now return to the 4 migrating models with $\varphi_{0}=$ 0.0099, 0.0056, 0.0042 and 0.0023. All of these models display irregular oscillations in the central value of the lapse function around some equilibrium position. The total integrated mass at the boundary as a function of time for these models is shown in Figure 15. Here, we again observe a fractional loss of the initial mass due to the fact that some of the field is ejected to infinity (in all these cases approximately 35-39%).
@@ V Discussion and conclusions
In this paper we have considered families of solutions for excited states of Proca stars in spherical symmetry, with a focus on exploring the dynamical behavior of the first two excited states under perturbations. We considered both the families of stationary solutions, as well as non-linear dynamical simulations of a set of self-consistent perturbed models. Our findings offer valuable insights into the stability and evolution of excited Proca stars, contributing to a further understanding of these compact objects.
It is well known that spherical Proca stars in the ground state are stable against small perturbations for a specific branch corresponding to those configurations with negative binding energy and mass to the left of the maximum mass (minimum binding energy) in the mass versus $\varphi_{0}$ diagram. However, we find that in the case of Proca stars in the first and second excited states this stable branch does not exist. In fact, we find that for excited states only a metastable region exists for configurations with negative binding energy and low masses. In this metastable branch the Proca stars migrate under perturbations to solutions with lower mass in the stable branch of the ground state, while ejecting excess field to infinity. The behavior of excited Proca stars is then very similar to the one found in the case of scalar boson stars, where dynamical evolutions have shown that excited states are unstable, and they either decay to the fundamental state or collapse to a black hole [30]. Our dynamical simulations of excited Proca stars have allowed us not only to observe the migration phenomena but also to determine the final configuration on the ground state to which they are migrating.
We should stress the fact that our models do not incorporate self-interactions. It has been shown that self-interactions can stabilize excited states of scalar boson stars [46, 47]. However, recent studies demonstrate that in the case of vector fields self-interactions can lead to significant pathologies, particularly the loss of hyperbolicity of the field equations [48, 49]. Introducing self-interaction parameters into the Proca star model therefore requires careful consideration of these instabilities, as highlighted by recent findings [50, 51, 52, 53]. This presents an opportunity to further explore the stability and dynamics of excited Proca stars, and to understand the complexities associated with self-interactions. We are currently in the process of studying this further.